# Distributionally and Adversarially Robust Logistic Regression via Intersecting Wasserstein Balls

### v1

@@ Appendix B Proofs
For any $\bm{\beta}\in\mathbb{R}^{n}$, with standard robust optimization arguments [4, 8], we can show that
@@
where the first step follows from the definition of logloss, the second step follows from the fact that $\log$ and $\exp$ are increasing functions, the third step takes the constant terms out of the $\sup$ problem and exploits the fact that the optimal solution of maximizing a linear function will be at an extreme point of the $\ell_{p}$ ball, the fourth step uses the definition of dual norm, and finally the redundant $-y\in\{-1,+1\}$ is omitted from the dual norm. We can therefore define the adversarial loss $\ell^{\alpha}_{\bm{\beta}}(\bm{x},y):=\log(1+\exp(-y\cdot\bm{\beta}^{\top}\bm{x}+\alpha\cdot\lVert\bm{\beta}\rVert_{p^{\star}}))$ where $\alpha$ models the strength of the adversary, $\bm{\beta}$ is the decision vector, and $(\bm{x},y)$ is an instance. Replacing ${\sup_{\bm{z}:\lVert\bm{z}\rVert_{p}\leq\alpha}}\{\ell_{\bm{\beta}}(\bm{x}+\bm{z},y)\}$ in 8 with $\ell^{\alpha}_{\bm{\beta}}(\bm{x},y)$ concludes the equivalence.
Furthermore, to see $\mathrm{Lip}(L^{\alpha})=1$, firstly note that since $L^{\alpha}(z)=\log(1+\exp(-z+\alpha\cdot\lVert\bm{\beta}\|_{p^{\star}}))$ is differentiable everywhere in $z$ and its gradient ${L^{\alpha}}^{\prime}$ is bounded everywhere, we have that $\mathrm{Lip}(L^{\alpha})$ is equal to $\sup_{z\in\mathbb{R}}\{|{L^{\alpha}}^{\prime}(z)|\}$. We thus have
@@
B.2 Proof of Corollary 4.3
Proposition 4.1 lets us represent DR-ARO as the DR counterpart of empirical minimization of $\ell^{\alpha}_{\bm{\beta}}$:
@@
Since the univariate loss $L^{\alpha}(z):=\log(1+\exp(-z+\alpha\cdot\lVert\bm{\beta}\rVert_{p^{\star}}))$ satisfying the identity $L^{\alpha}(\langle y\cdot\bm{x},\bm{\beta}\rangle)=\ell^{\alpha}_{\bm{\beta}}(\bm{x},y)$ is Lipschitz continuous (cf. Proposition 4.1), Theorem 14 (ii) of [59] is immediately applicable. We can therefore rewrite (19) as:
@@
B.3 Proof of Proposition 5.2
We prove Proposition 5.2 by constructing the optimization problem in its statement. We will thus dualize the inner $\sup$ problem of Inter-ARO for fixed $\bm{\beta}$. To this end, we present a sequence of reformulations to the inner problem and then exploit strong semi-infinite duality.
@@
The following semi-infinite optimization problem, obtained by standard algebraic duality, is a strong dual to the above problem since $\varepsilon,\widehat{\varepsilon}>0$ [61].
@@
B.4 Proof of Proposition 5.3
We first present a technical lemma that will allow us to rewrite a specific type of difference of convex functions (DC) maximization problem that appears in the constraints of Inter-ARO. Rewriting such DC maximization problems is one of the key steps in reformulating Wasserstein DRO problems, and our lemma is inspired from [59, Lemma 47][58, Theorem 3.8], and [3, Lemma 1] who reformulate maximizing the difference of a convex function and a norm. Our DRO problem Inter-ARO, however, comprises two ambiguity sets, hence the DC term that we investigate will be the difference between a convex function and the sum of two norms. This requires a new analysis and we will see that Inter-ARO is NP-hard due to this additional difficulty.
Lemma B.1.
@@
where the first identity follows from the definition of composition and the second identity employs Toland’s duality [69] to rewrite difference of convex functions optimization.
By using infimal convolutions [52, Theorem 16.4], we can reformulate $g^{*}$:
@@
DC maximization terms similar to the one dealt by Lemma B.1 appear on the left-hand side of the constraints of Inter-ARO (cf. formulation in Proposition 5.2). These constraints would admit a tractable reformulation for the case without auxiliary data because the $\inf$ term in the reformulation presented in Lemma B.1 does not appear in such cases. To see this, eliminate the second norm (the one associated with auxiliary data) by taking $\widehat{\lambda}=0$, which will cause the constraint $\lvert\theta\rvert\cdot\lVert\bm{z}\rVert_{q^{\star}}\leq\widehat{\lambda}$ to force $\bm{z}=\mathbf{0}$, and the alternative formulation will thus be:
@@
where we used the fact that $L=L^{**}$ and $\sup_{\theta\in\mathrm{dom}(L)}\lvert\theta\rvert=\mathrm{Lip}(L)$ since $L$ is closed convex [52, Corollary 13.3.3]. Hence, the DC maximization can be represented with a convex function with an additional convex inequality, making the constraints tractable for the case without auxiliary data. For the case with auxiliary data, however, the $\sup_{\theta}\inf_{\bm{z}}$ structure makes these constraints equivalent to two-stage robust constraints (with uncertain parameter $\theta$ and adjustable variable $\bm{z}$), bringing an adjustable robust optimization ([5, 77]) perspective to Inter-ARO. By using the univariate representation $\ell_{\bm{\beta}}^{\alpha}(\bm{x},y)=L^{\alpha}(y\cdot\bm{\beta}^{\top}\bm{x})$Inter-ARO can be written as
@@
and applying Lemma B.1 to the left-hand side of the constraints gives:
@@
By using adjustable robust optimization theory, we show that this problem is NP-hard even in the simplest setting. To this end, take $N=\widehat{N}=1$ as well as $\kappa=0$; the formulation presented in Proposition 5.2 reduces to:
@@
and an application of Lemma B.1 leads us to the following reformulation:
@@
This reduces problem (36) to a generic two-stage robust optimization problem with random recourse [66, Problem 1] which is proven to be NP-hard even if $S^{\alpha*}$ was constant [31]. ∎
B.5 Proof of Theorem 5.4
Consider the reformulation 33 of Inter-ARO that we introduced in the proof of Proposition 5.3. For any $i\in[N]$ and $j\in[\widehat{N}]$, the corresponding constraint in the first group of ‘adjustable robust’ ($\forall$,\ $\exists$ , ∃) constraints will be:
@@
Notice that this is a safe approximation, since any fixed $\bm{z}$ satisfying the latter system is a feasible static solution in the former system, meaning that for every realization of $\theta$ in the first system, the inner $\exists\bm{z}$ can always ‘play’ the same $\bm{z}$ that is feasible in the latter system (hence the latter is named the static relaxation, [10]). In the relaxed system, we can drop $\forall\theta$ and keep its worst-case realization instead:
@@
The term $\sup_{\theta\in\mathrm{dom}(L^{*})}\{-L^{\alpha*}(\theta)+\theta\cdot\bm{\beta}^{\top}\bm{x}^{i}+\theta\cdot\bm{z}^{\top}(\widehat{\bm{x}}^{j}-\bm{x}^{i})\}$ is the definition of the biconjugate $L^{\alpha**}(\bm{\beta}^{\top}\bm{x}^{i}+\bm{z}^{\top}(\widehat{\bm{x}}^{j}-\bm{x}^{i}))$. Since $L^{\alpha}$ is a closed convex function, we have $L^{\alpha**}=L^{\alpha}$ [52, Corollary 12.2.1]. Moreover, $\sup_{\theta\in\mathrm{dom}(L^{*})}\{\lvert\theta\rvert\}$ is an alternative representation of the Lipschitz constant of the function $L^{\alpha}$ [52, Corollary 13.3.3], which is equal to $1$ as we showed earlier. The adjustable robust constraint thus reduces to:
@@
as a result of the static relaxation. This relaxed reformulation applies to all $i\in[N]$ and $j\in[\widehat{N}]$ as well as to the second group of adjustable robust constraints analogously. Replacing each constraint of 33 with this system concludes the proof.
B.6 Proof of Corollary 5.6
@@
where we also used the definition of $L^{\alpha}$. This problem is identical to DR-ARO, which means that feasible solutions of DR-ARO are feasible for 13 if the additional variables $(\widehat{\lambda},\widehat{\bm{s}},\bm{z}^{l}_{ij})$ are set to zero, concluding the first statement of the corollary.
The second statement is immediate since $\widehat{\varepsilon}\rightarrow\infty$ forces $\widehat{\lambda}=0$ due to the term $\widehat{\varepsilon}\widehat{\lambda}$ in the objective of 13, and this proof shows in such a case 13 reduces to DR-ARO (which is identical to Inter-ARO when $\varepsilon\rightarrow\infty$ by definition).
B.7 Proof of Proposition 5.7
@@
which coincides with the objective function of (14). The proof of Proposition 4.1 shows
@@
which concludes the proof.
B.8 Proof of Proposition 5.8
@@
and since $d$ is a feature-label metric (cf. Definition 3.1) the minimum is well-defined [73, Theorem 4.1]. We name the optimal solutions to the above problem the optimal couplings. Let $\Pi^{1}$ be an optimal coupling of $\mathrm{W}(\mathbb{Q},\mathbb{P}^{1})$ and let $\Pi^{2}$ be an optimal coupling of $\mathrm{W}(\mathbb{Q},\mathbb{P}^{2})$ and define $\Pi^{c}=\lambda\cdot\Pi^{1}+(1-\lambda)\cdot\Pi^{2}$. We have
@@
where the first identity uses the definition of the Wasserstein metric, the inequality is due to Lemma B.2 as $\Pi^{c}$ is a feasible coupling (not necessarily optimal), the equality that follows uses the definition of $\Pi^{c}$ and the linearity of integrals, and the final identity uses the fact that $\Pi^{1}$ and $\Pi^{2}$ were constructed to be the optimal couplings. ∎
We now prove the proposition (we refer to $\mathbb{Q}_{\mathrm{mix}}$ in the statement of this lemma simply as $\mathbb{Q}$). To prove $\mathbb{Q}\in\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})\cap\mathfrak{B}_{\widehat{\varepsilon}}({\widehat{\mathbb{P}}}_{\widehat{N}})$, it is sufficient to show that $\mathrm{W}(\mathbb{P}_{N},\mathbb{Q})\leq\varepsilon$ and $\mathrm{W}(\widehat{\mathbb{P}}_{\widehat{N}},\mathbb{Q})\leq\widehat{\varepsilon}$ jointly hold. By using Lemma B.3, we can derive the following inequalities:
@@
which is implied when $\dfrac{\lambda}{1-\lambda}=\dfrac{\widehat{\varepsilon}}{\varepsilon}$, concluding the proof.
B.9 Proof of Theorem 6.1
Since each result in the statement of this theorem is abridged, we will present these results sequentially as separate results. We review the existing literature to characterize $\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})$, in a similar fashion with the results presented in [55, Appendix A] for the logistic loss, by revising them to the adversarial loss whenever necessary. The $N$-fold product distribution of $\mathbb{P}^{0}$ from which the training set $\mathbb{P}_{N}$ is constructed is denoted below by $[\mathbb{P}^{0}]^{N}$.
Theorem B.4.
@@
The statement follows from Theorem 18 of [34]. The presented decay rate $\mathcal{O}(N^{-1/n})$ of $\varepsilon$ as $N$ increases is optimal [25]. ∎
Now that we gave a confidence for the radius $\varepsilon$ of $\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})$, we analyze the underlying optimization problems. Most of the theory is well-established for logistic loss function, and in the following we show that similar results follow for the adversarial loss function. For convenience, we state DR-ARO again by using the adversarial loss function as in the proof of Proposition 4.1:
@@
Theorem B.5.
If the assumptions of Theorem B.4 are satisfied and $\varepsilon$ is chosen as in the statement of Theorem B.4, then
@@
holds for all $\eta\in(0,1)$ and all optimizers $\bm{\beta}^{\star}$ of 39.
@@
The statement follows from Theorem 19 of [34] given that $\ell^{\alpha}_{\bm{\beta}}$ is a finite-valued continuous loss function. ∎
Theorem B.5 states that the optimal value of 39 overestimates the true loss with arbitrarily high confidence $1-\eta$. Despite the desired overestimation of the true loss, we show that 39 is still asymptotically consistent if we restrict the set of admissible $\bm{\beta}$ to a bounded set [^1] Note that, this is without loss of generality given that we can normalize the decision boundary of linear classifiers..
Theorem B.6.
If we restrict the hypotheses $\bm{\beta}$ to a bounded set $\mathcal{H}\subseteq\mathbb{R}^{n}$, and parameterize $\varepsilon$ as $\varepsilon_{N}$ to show its dependency to the sample size, then, under the assumptions of Theorem B.4, we have
@@
whenever $\varepsilon_{N}$ is set as specified in Theorem B.4 along with its finite-sample confidence $\eta_{N}$, and they satisfy $\sum_{N\in\mathbb{N}}\eta_{N}<\infty$ and $\lim_{N\rightarrow\infty}\varepsilon_{N}=0$.
@@
If we show that there exists $\bm{\xi}^{0}\in\Xi$ and $C>0$ such that $\ell^{\alpha}_{\bm{\beta}}(\bm{x},y)\leq C(1+d(\bm{\xi},\bm{\xi}^{0}))$ holds for all $\bm{\beta}\in\mathcal{H}$ and $\bm{\xi}\in\Xi$ (that is, the adversarial loss satisfies a growth condition), the statement will follow immediately from Theorem 20 of [34].
@@
$\log(1+\exp(-y\cdot\bm{\beta}^{\top}\bm{x}+\alpha\cdot\lVert\bm{\beta}\rVert_{p^{\star}}))\leq C(1+\lVert\bm{x}-\bm{x}^{0}\rVert_{q}+\kappa\cdot\mathds{1}\small{[y\neq y^{0}]})$
@@
$C(1+\lVert\bm{x}-\bm{x}^{0}\rVert_{q}+\kappa\cdot\mathds{1}\small{[y\neq y^{0}]})$ $=C(1+\lVert\bm{x}\rVert_{q}+\kappa\cdot\mathds{1}\small{[y\neq y^{0}]})$
@@
So far, we reviewed tight characterizations for $\varepsilon$ so that the ball $\mathfrak{B}_{\varepsilon}({{\mathbb{P}}}_{{N}})$ includes the true distribution $\mathbb{P}^{0}$ with arbitrarily high confidence, proved that the DRO problem 39 overestimates the true loss, while converging to the true problem asymptotically as the confidence $1-\eta$ increases and the radius $\varepsilon$ decreases simultaneously. Finally, we discuss that for optimal solutions $\bm{\beta}^{\star}$ to 39, there are worst case distributions $\mathbb{Q}^{\star}\in\mathfrak{B}_{\varepsilon}({{\mathbb{P}}}_{{N}})$ of nature’s problem that are supported on at most $N+1$ atoms.
Theorem B.7.
@@
The proof follows from [80]. ∎
See the proof of [55, Theorem 8] and the discussion that follows for insights and further analysis on these results presented.
B.10 Proof of Theorem 6.2
Firstly, since ^$\widehat{\mathbb{P}}_{\widehat{N}}$ is constructed from i.i.d. samples of $\widehat{\mathbb{P}}$, we can overestimate the distance $\widehat{\varepsilon}_{1}=\mathrm{W}(\widehat{\mathbb{P}}_{\widehat{N}},\widehat{\mathbb{P}})$ analogously by applying Theorem B.4, mutatis mutandis. This leads us to the following result where the joint (independent) $N$-fold product distribution of $\mathbb{P}^{0}$ and the $\widehat{N}$-fold product distribution of $\widehat{\mathbb{P}}$ is denoted below by $[\mathbb{P}^{0}\times\widehat{\mathbb{P}}]^{N\times\widehat{N}}$.
Theorem B.8.

### v2

@@ 10.1 Proof of Observation 1
For any $\bm{\beta}\in\mathbb{R}^{n}$, with standard robust optimization arguments (Ben-Tal et al.,, 2009; Bertsimas and Den Hertog,, 2022), we can show that
@@
where the first step follows from the definition of logloss, the second step follows from the fact that $\log$ and $\exp$ are increasing functions, the third step takes the constant terms out of the $\sup$ problem and exploits the fact that the optimal solution of maximizing a linear function will be at an extreme point of the $\ell_{p}$ ball, the fourth step uses the definition of dual norm, and finally the redundant $-y\in\{-1,+1\}$ is omitted from the dual norm. We can therefore define the adversarial loss $\ell^{\alpha}_{\bm{\beta}}(\bm{x},y):=\log(1+\exp(-y\cdot\bm{\beta}^{\top}\bm{x}+\alpha\cdot\lVert\bm{\beta}\rVert_{p^{\star}}))$ where $\alpha$ models the strength of the adversary, $\bm{\beta}$ is the decision vector, and $(\bm{x},y)$ is an instance. Replacing ${\sup_{\bm{z}:\lVert\bm{z}\rVert_{p}\leq\alpha}}\{\ell_{\bm{\beta}}(\bm{x}+\bm{z},y)\}$ in DR-ARO with $\ell^{\alpha}_{\bm{\beta}}(\bm{x},y)$ concludes the equivalence of the optimization problem.
Furthermore, to see $\mathrm{Lip}(L^{\alpha})=1$, firstly note that since $L^{\alpha}(z)=\log(1+\exp(-z+\alpha\cdot\lVert\bm{\beta}\|_{p^{\star}}))$ is differentiable everywhere in $z$ and its gradient ${L^{\alpha}}^{\prime}$ is bounded everywhere, we have that $\mathrm{Lip}(L^{\alpha})$ is equal to $\sup_{z\in\mathbb{R}}\{|{L^{\alpha}}^{\prime}(z)|\}$. We thus have:
@@
10.2 Proof of Corollary 1
Observation 1 lets us represent DR-ARO as the DR counterpart of empirical minimization of $\ell^{\alpha}_{\bm{\beta}}$:
@@ 10.2 Proof of Corollary 1
Since the univariate loss $L^{\alpha}(z):=\log(1+\exp(-z+\alpha\cdot\lVert\bm{\beta}\rVert_{p^{\star}}))$ satisfying the identity $L^{\alpha}(\langle y\cdot\bm{x},\bm{\beta}\rangle)=\ell^{\alpha}_{\bm{\beta}}(\bm{x},y)$ is Lipschitz continuous, Theorem 14 (ii) of Shafieezadeh-Abadeh et al., (2019) is immediately applicable. We can therefore rewrite (24) as:
@@
10.3 Proof of Proposition 1
We prove Proposition 1 by constructing the optimization problem in its statement. We will thus dualize the inner $\sup$ problem of Inter-ARO for fixed $\bm{\beta}$. To this end, we present a sequence of reformulations to the inner problem and then exploit strong semi-infinite duality.
@@ 10.3 Proof of Proposition 1
The following semi-infinite optimization problem, obtained by standard algebraic duality, is a strong dual to the above problem since $\varepsilon,\widehat{\varepsilon}>0$ (Shapiro,, 2001).
@@
10.4 Proof of Proposition 2
We first present a technical lemma that will allow us to rewrite a specific type of difference of convex functions (DC) maximization problem that appears in the constraints of Inter-ARO. Rewriting such DC maximization problems is one of the key steps in reformulating Wasserstein DRO problems, and our lemma is inspired from Shafieezadeh-Abadeh et al., (2019, Lemma 47)Shafieezadeh-Abadeh et al., (2023, Theorem 3.8), and Belbasi et al., (2023, Lemma 1) who reformulate maximizing the difference of a convex function and a norm. Our DRO problem Inter-ARO, however, comprises two ambiguity sets, hence the DC term that we investigate will be the difference between a convex function and the sum of two norms. This requires a new analysis and we will see that Inter-ARO is NP-hard due to this additional difficulty.
Lemma 1.
@@ 10.4 Proof of Proposition 2
where the first identity follows from the definition of composition and the second identity employs Toland’s duality (Toland,, 1978) to rewrite difference of convex functions optimization.
By using infimal convolutions (Rockafellar,, 1997, Theorem 16.4), we can reformulate $g^{*}$:
@@
DC maximization terms similar to the one dealt by Lemma 1 appear on the left-hand side of the constraints of Inter-ARO (cf. formulation in Proposition 1). These constraints would admit a tractable reformulation for the case without auxiliary data because the $\inf$ term in the reformulation presented in Lemma 1 does not appear in such cases. To see this, eliminate the second norm (the one associated with auxiliary data) by taking $\widehat{\lambda}=0$, which will cause the constraint $\lvert\theta\rvert\cdot\lVert\bm{z}\rVert_{q^{\star}}\leq\widehat{\lambda}$ to force $\bm{z}=\mathbf{0}$, and the alternative formulation will thus be:
@@
where we used the fact that $L=L^{**}$ and $\sup_{\theta\in\mathrm{dom}(L)}\lvert\theta\rvert=\mathrm{Lip}(L)$ since $L$ is closed convex (Rockafellar,, 1997, Corollary 13.3.3). Hence, the DC maximization can be represented with a convex function with an additional convex inequality, making the constraints tractable for the case without auxiliary data. For the case with auxiliary data, however, the $\sup_{\theta}\inf_{\bm{z}}$ structure makes these constraints equivalent to two-stage robust constraints (with uncertain parameter $\theta$ and adjustable variable $\bm{z}$), bringing an adjustable robust optimization (Ben-Tal et al.,, 2004; Yanıkoğlu et al.,, 2019) perspective to Inter-ARO. By using the univariate representation $\ell_{\bm{\beta}}^{\alpha}(\bm{x},y)=L^{\alpha}(y\cdot\bm{\beta}^{\top}\bm{x})$Inter-ARO can be written as
@@
and applying Lemma 1 to the left-hand side of the constraints gives:
@@
By using adjustable robust optimization theory, we show that this problem is NP-hard even in the simplest setting. To this end, take $N=\widehat{N}=1$ as well as $\kappa=0$; the formulation presented in Proposition 1 reduces to:
@@
and an application of Lemma 1 leads us to the following reformulation:
@@
This reduces problem (41) to a generic two-stage robust optimization problem with random recourse (Subramanyam et al.,, 2020, Problem 1) which is proven to be NP-hard even if $S^{\alpha*}$ was constant (Guslitser,, 2002). ∎
10.5 Proof of Theorem 1
Consider the reformulation 38 of Inter-ARO that we introduced in the proof of Proposition 2. For any $i\in[N]$ and $j\in[\widehat{N}]$, the corresponding constraint in the first group of ‘adjustable robust’ ($\forall$,\ $\exists$ , ∃) constraints will be:
@@ 10.5 Proof of Theorem 1
Notice that this is a safe approximation, since any fixed $\bm{z}$ satisfying the latter system is a feasible static solution in the former system, meaning that for every realization of $\theta$ in the first system, the inner $\exists\bm{z}$ can always ‘play’ the same $\bm{z}$ that is feasible in the latter system (hence the latter is named the static relaxation, Bertsimas et al., 2015). In the relaxed system, we can drop $\forall\theta$ and keep its worst-case realization instead:
@@
The term $\sup_{\theta\in\mathrm{dom}(L^{*})}\{-L^{\alpha*}(\theta)+\theta\cdot\bm{\beta}^{\top}\bm{x}^{i}+\theta\cdot\bm{z}^{\top}(\widehat{\bm{x}}^{j}-\bm{x}^{i})\}$ is the definition of the biconjugate $L^{\alpha**}(\bm{\beta}^{\top}\bm{x}^{i}+\bm{z}^{\top}(\widehat{\bm{x}}^{j}-\bm{x}^{i}))$. Since $L^{\alpha}$ is a closed convex function, we have $L^{\alpha**}=L^{\alpha}$ (Rockafellar,, 1997, Corollary 12.2.1). Moreover, $\sup_{\theta\in\mathrm{dom}(L^{*})}\{\lvert\theta\rvert\}$ is an alternative representation of the Lipschitz constant of the function $L^{\alpha}$ (Rockafellar,, 1997, Corollary 13.3.3), which is equal to $1$ as we showed earlier. The adjustable robust constraint thus reduces to:
@@
as a result of the static relaxation. This relaxed reformulation applies to all $i\in[N]$ and $j\in[\widehat{N}]$ as well as to the second group of adjustable robust constraints analogously. Replacing each constraint of 38 with this system concludes the proof.∎
10.6 Proof of Corollary 2
@@
where we also used the definition of $L^{\alpha}$. This problem is identical to DR-ARO, which means that feasible solutions of DR-ARO are feasible for 16 if the additional variables $(\widehat{\lambda},\widehat{\bm{s}},\bm{z}^{l}_{ij})$ are set to zero, concluding the first statement of the corollary.
The second statement is immediate since $\widehat{\varepsilon}\rightarrow\infty$ forces $\widehat{\lambda}=0$ due to the term $\widehat{\varepsilon}\widehat{\lambda}$ in the objective of 16, and this proof shows in such a case 16 reduces to DR-ARO (which is identical to Inter-ARO when $\varepsilon\rightarrow\infty$ by definition). ∎
10.7 Proof of Observation 2
@@
which coincides with the objective function of (19). Since we have
@@
we can conclude the proof. ∎
10.8 Proof of Proposition 3
@@
and since $d$ is a feature-label metric (cf. Definition 1) the minimum is well-defined (Villani et al.,, 2009, Theorem 4.1). We name the optimal solutions to the above problem the optimal couplings. Let $\Pi^{1}$ be an optimal coupling of $\mathrm{W}(\mathbb{Q},\mathbb{P}^{1})$ and let $\Pi^{2}$ be an optimal coupling of $\mathrm{W}(\mathbb{Q},\mathbb{P}^{2})$ and define $\Pi^{c}=\lambda\cdot\Pi^{1}+(1-\lambda)\cdot\Pi^{2}$. We have
@@
where the first identity uses the definition of the Wasserstein metric, the inequality is due to Lemma 2 as $\Pi^{c}$ is a feasible coupling (not necessarily optimal), the equality that follows uses the definition of $\Pi^{c}$ and the linearity of integrals, and the final identity uses the fact that $\Pi^{1}$ and $\Pi^{2}$ were constructed to be the optimal couplings. ∎
We now prove the proposition (we refer to $\mathbb{Q}_{\mathrm{mix}}$ in the statement of this lemma simply as $\mathbb{Q}$). To prove $\mathbb{Q}\in\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})\cap\mathfrak{B}_{\widehat{\varepsilon}}({\widehat{\mathbb{P}}}_{\widehat{N}})$, it is sufficient to show that $\mathrm{W}(\mathbb{P}_{N},\mathbb{Q})\leq\varepsilon$ and $\mathrm{W}(\widehat{\mathbb{P}}_{\widehat{N}},\mathbb{Q})\leq\widehat{\varepsilon}$ jointly hold. By using Lemma 3, we can derive the following inequalities:
@@
which is implied when $\dfrac{\lambda}{1-\lambda}=\dfrac{\widehat{\varepsilon}}{\varepsilon}$, concluding the proof. ∎
10.9 Proof of Theorem 2
Since each result in the statement of this theorem is abridged, we will present these results sequentially as separate results. We review the existing literature to characterize $\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})$, in a similar fashion with the results presented in (Selvi et al.,, 2022, Appendix A) for the logistic loss, by revising them to the adversarial loss whenever necessary. The $N$-fold product distribution of $\mathbb{P}^{0}$ from which the training set $\mathbb{P}_{N}$ is constructed is denoted below by $[\mathbb{P}^{0}]^{N}$.
Theorem 4.
@@ 10.9 Proof of Theorem 2
The statement follows from Theorem 18 of Kuhn et al., (2019). The presented decay rate $\mathcal{O}(N^{-1/n})$ of $\varepsilon$ as $N$ increases is optimal (Fournier and Guillin,, 2015). ∎
Now that we gave a confidence for the radius $\varepsilon$ of $\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})$, we analyze the underlying optimization problems. Most of the theory is well-established for logistic loss function, and in the following we show that similar results follow for the adversarial loss function. For convenience, we state DR-ARO again by using the adversarial loss function as defined in Observation 1:
@@
Theorem 5.
If the assumptions of Theorem 4 are satisfied and $\varepsilon$ is chosen as in the statement of Theorem 4, then
@@
holds for all $\eta\in(0,1)$ and all optimizers $\bm{\beta}^{\star}$ of 44.
@@
The statement follows from Theorem 19 of Kuhn et al., (2019) given that $\ell^{\alpha}_{\bm{\beta}}$ is a finite-valued continuous loss function. ∎
Theorem 5 states that the optimal value of 44 overestimates the true loss with arbitrarily high confidence $1-\eta$. Despite the desired overestimation of the true loss, we show that 44 is still asymptotically consistent if we restrict the set of admissible $\bm{\beta}$ to a bounded set [^2] Note that, this is without loss of generality given that we can normalize the decision boundary of linear classifiers..
Theorem 6.
If we restrict the hypotheses $\bm{\beta}$ to a bounded set $\mathcal{H}\subseteq\mathbb{R}^{n}$, and parameterize $\varepsilon$ as $\varepsilon_{N}$ to show its dependency to the sample size, then, under the assumptions of Theorem 4, we have
@@
whenever $\varepsilon_{N}$ is set as specified in Theorem 4 along with its finite-sample confidence $\eta_{N}$, and they satisfy $\sum_{N\in\mathbb{N}}\eta_{N}<\infty$ and $\lim_{N\rightarrow\infty}\varepsilon_{N}=0$.
@@
If we show that there exists $\bm{\xi}^{0}\in\Xi$ and $C>0$ such that $\ell^{\alpha}_{\bm{\beta}}(\bm{x},y)\leq C(1+d(\bm{\xi},\bm{\xi}^{0}))$ holds for all $\bm{\beta}\in\mathcal{H}$ and $\bm{\xi}\in\Xi$ (that is, the adversarial loss satisfies a growth condition), the statement will follow immediately from Theorem 20 of (Kuhn et al.,, 2019).
@@
$\log(1+\exp(-y\cdot\bm{\beta}^{\top}\bm{x}+\alpha\cdot\lVert\bm{\beta}\rVert_{p^{\star}}))\leq C(1+\lVert\bm{x}-\bm{x}^{0}\rVert_{q}+\kappa\cdot\mathds{1}{[y\neq y^{0}]})$
@@
$C(1+\lVert\bm{x}-\bm{x}^{0}\rVert_{q}+\kappa\cdot\mathds{1}{[y\neq y^{0}]})$ $=C(1+\lVert\bm{x}\rVert_{q}+\kappa\cdot\mathds{1}{[y\neq y^{0}]})$
@@
So far, we reviewed tight characterizations for $\varepsilon$ so that the ball $\mathfrak{B}_{\varepsilon}({{\mathbb{P}}}_{{N}})$ includes the true distribution $\mathbb{P}^{0}$ with arbitrarily high confidence, proved that the DRO problem 44 overestimates the true loss, while converging to the true problem asymptotically as the confidence $1-\eta$ increases and the radius $\varepsilon$ decreases simultaneously. Finally, we discuss that for optimal solutions $\bm{\beta}^{\star}$ to 44, there are worst case distributions $\mathbb{Q}^{\star}\in\mathfrak{B}_{\varepsilon}({{\mathbb{P}}}_{{N}})$ of nature’s problem that are supported on at most $N+1$ atoms.
Theorem 7.
@@
The proof follows from (Yue et al.,, 2022). ∎
See the proof of Selvi et al., (2022, Theorem 8) and the discussion that follows for insights and further analysis on these results presented.
10.10 Proof of Theorem 3
Firstly, since ^$\widehat{\mathbb{P}}_{\widehat{N}}$ is constructed from i.i.d. samples of $\widehat{\mathbb{P}}$, we can overestimate the distance $\widehat{\varepsilon}_{1}=\mathrm{W}(\widehat{\mathbb{P}}_{\widehat{N}},\widehat{\mathbb{P}})$ analogously by applying Theorem 4, mutatis mutandis. This leads us to the following result where the joint (independent) $N$-fold product distribution of $\mathbb{P}^{0}$ and the $\widehat{N}$-fold product distribution of $\widehat{\mathbb{P}}$ is denoted below by $[\mathbb{P}^{0}\times\widehat{\mathbb{P}}]^{N\times\widehat{N}}$.
Theorem 8.